Find the equivalent circuit with fewest elements of the following network.Question 2b. (15 marks) Find the equivalent circuit with the fewest elements of the following network. (15 marks) 6 mH 000 10
Answer #1

The  10 \ \mu F and  90 \ \mu F are in series, their effective series capacitance is

\\C_s = \frac{10 * 90}{10 + 90} \\\\C_s = 9 \ \mu F

Now, this  \\C_s is in parallel with the  9 \ \mu F capacitor

therefore, the total capacitance is

\\C_{total} = C_s + 9 \\\\C_{total} = 9 + 9 \\\\C_{total} = 18 \ \mu C

The two inductors 14 \ mH and 35 \ mH are in parallel, their effective parallel inductance is

\\L_p = \frac{14 * 35}{14 + 35} \\\\L_p = 10 \ mH

Now this  L_p is in series with the  6 \ mH inductor,

therefore, the total inductance is,

\\L_{total} = L_p + 6 \\\\L_{total} = 10 + 6 \\\\L_{total} = 16 \ mH

Therefore, the equivalent circuit with the fewest elements is

L total 16 mol etotal 18 iC